Luke's Lubricants starts business on January 1. The following operations data are available for January for the one lubricanti produces. 


Gallons Beginning inventory 0

Started in January 160,000

Ending work-in-process inventory (70% complete) 15,000 


Costs incurred in January follow. 

Materials $ 167, 125

Labor 28,000

Manufacturing overhead 77,000


All production at Luke's is sold as it is produced there are no finished goods inventories). 


Required: 

a. Compute cost of goods sold for January. 

b. What is the value of work in process inventory on January 31?
Answer #4

=>160,000-15,000=14,500

=>15,000*70%=10,500

=>14,500+10,500=155,500

=>167,125+28,000+77,000=272,125

=>272,125/155,500=1.75

=>14,500*1.75=253,750=>answer for part A

=>10,500*1.75=18,375=>answer for part B

answered by: xyzlmnop
Answer #1

Unit transferred out = 160000-15000 = 145000

Equivalent unit =145000+(15000*70%) = 155500

Cost per equivalent unit = 272125/155500 = 1.75

a) Cost of goods sold = 145000*1.75 = 253750

b) Work in process inventory on January 31 = 10500*1.75 = 18375
Answer #2

from costs incurred 167,125+ 28,000+ 77,000= 272,125
Answer #3

15,000* 70%= 10,500 for part b